The manager of a restaurant found that the cost to produce 50 cups of coffee is
$ 9.00​, while the cost to produce 350 cups is ​$39.00. Assume the relationship between the cost y to produce x cups of coffee is linear.
a. Write a linear equation that expresses the​ cost, y, in terms of the number of cups of​ coffee, x.
b.  How many cups of coffee are produced if the cost of production is ​$

Answer by math_tutor2020(3817)


x = number of cups of coffee
y = total cost in dollars

50 cups cost $9, meaning: x = 50 and y = 9
I.e. the point (x,y) = (50, 9) is on the line

Another point is (350, 39) since 350 cups cost $39 to produce.

Let's find the slope of the line through (50,9) and (350,39)
(x1,y1) = (50,9) and (x2,y2) = (350,39)
m = (y2 - y1)/(x2 - x1)
m = (39 - 9)/(350 - 50)
m = 30/300
m = 1/10
m = 0.10

The slope of 0.10 means that each time x goes up by 1, y increases by 0.10
Each new cup of coffee costs $0.10, i.e. 10 cents

Now apply point slope form.
y - y1 = m(x - x1)
y - 9 = 0.10(x - 50)
y - 9 = 0.10x - 0.10*50
y - 9 = 0.10x - 5
y = 0.10x - 5+9
y = 0.10x + 4

The equation is in y = mx+b form
m = 0.10 = slope
b = 4 = y intercept

The slope tells us how much each extra cup costs
The y intercept is the starting cost. If zero cups are produced, there is still a fixed fee of $4 to set everything up. Think of it like a fee to get in the door.

All of the work/steps shown above apply to part (a) only.

----------------------------------

Part (b) cannot be answered since the cost of production has been cut off.
You'll need to revise your post to include this value.

I'll go over a hypothetical example
Let's say the cost of production is $100

This means we plug in y = 100 and solve for x.
y = 0.10x + 4
0.10x + 4 = y
0.10x + 4 = 100
0.10x = 100-4
0.10x = 96
x = 96/(0.10)
x = 960
Therefore, a production cost of $100 means that 960 cups were produced.

Once again, this is a hypothetical example. Follow steps similar to this with the actual cost your teacher gave you.